find and an equation equation of the line that satisfies the given conditions through (-4,4): perpendicular to the line y=2/3x-2
Answer by stanbon(75887)

find and an equation equation of the line that satisfies the given conditions through (-4,4): perpendicular to the line y=2/3x-2
----
The given equation has slope = 2/3
---
Any perpendicular line must have slope = -3/2
------
Form:: y = mx + b
Solve for "b"::
4 = (-3/2)(-4) + b
---------------
4 = 6 + b
b = -2
-------------
Equation:
y = (-3/2) - 2
-------------
Cheers,
Stan H.
--------------------